determine the Celsius temperature of 2.49 mol of gas contained in a 1.00-L vessel at a pressure of 143 kPa
1
Expert's answer
2020-04-10T13:38:09-0400

According to the equation of Mendeleev-Clapeyron for gases ("p•V = \nu•R•T"), we can determine the temperature using this formula (considering that Pa = N/m2 , L = m3 and J = N•m):

"T = (p•V) / (\nu•R)" = (143(N/m2) • 103 • 1(m3)) / (2.49(mol) • 8.31((N•m)/(mol•K))) = (143•103(N•m))/20.69((N•m)/K) = 6910 K = 6637oC

Answer: T=6637oC